# Individual Identification Using Radar-Measured Respiratory and Heartbeat Features

### v1

@@ Abstract
Index Terms:
classification, heartbeat features, individual identification, millimeter-wave radar, respiratory features
@@ I Introduction
Recently, radar has been used widely for measurement of human bodies because both limb movements and physiological signals can be acquired in a noncontact manner and privacy concerns are reduced when compared with use of camera-based sensors [12, 13, 14]. Radar techniques have been used for respiration measurement  [15, 16, 17], heartbeat [18, 19, 21, 20] and also for individual identification applications [22, 23].
@@
In this paper, we propose an individual identification method that uses both respiratory and heartbeat features for the first time. The contributions of this paper are listed as follows:
@@ II-A Waveform Feature Extraction Method
where $s(t)$ is the signal reflected from the target, $\mathrm{unwrap}\{\cdot\}$ is the phase unwrapping operator, and $\angle$ represents the phase of a complex number. When the target person is stationary, $d(t)$ mainly contains the displacement components that result from the person’s respiration and heartbeat.
@@
Example MRCW models are shown in Fig. 2 for $f=1$ Hz and $A=1$. As the figure shows, the waveform for the MRCW model is dependent on the parameters $\quantity(f,\beta_{1},\beta_{2},D)$. Hereafter, the parameters $\bm{p}=\quantity(f,\beta_{1},\beta_{2},D)^{\mathrm{T}}$ are used as the respiratory features.
@@
where $\tau$ is the time shift and $\hat{d}_{\mathrm{MRCW}}(t)$ is the MRCW model that provides the best approximation of $d(t)$. In addition, $w(t)$ is a rectangular window with a length of 8 s. The parameters estimated in (5) are defined here as the estimated waveform features $\hat{\bm{p}}(t)=(\hat{f}(t),\hat{\beta}_{1}(t),\hat{\beta}_{2}(t),\hat{D}(t))^{\mathrm{T}}$ at a time $t$.
@@ II-B Features of the Respiratory Expiratory Plateau
$w_{\mathrm{E}}^{(t)}(t^{\prime})=\begin{cases}w(t^{\prime}-t)&\quantity(\hat{d}_{\mathrm{MRCW}}(t^{\prime})\geq\varepsilon\hat{A}(t)),\\ 0&(\text{otherwise}),\end{cases}$ (8)
where $\varepsilon=0.6$ is a threshold that was set empirically. In this paper, $\hat{c}_{2}(t)$ is used as a feature that represents the expiratory plateau’s shape. Fig. 4 shows the time interval for the expiratory plateau (gray area) and the corresponding estimated parabolic curve (red solid line).
@@ III-A High-Frequency Enhancement Using Differentiation
In this section, we propose a method to extract the features of the heartbeat component required for individual identification. Because the fundamental frequency component of the heartbeat is often masked by harmonic components of the respiration, it has proven to be advantageous to use the higher harmonics of the heartbeat components [18]. Therefore, our proposed method emphasizes the high-frequency components by using second-order differentiation, as in our previous work [21]. Because the second-order differentiation $s^{\prime\prime}(t)=({\mathrm{d}}^{2}/{\mathrm{d}}t^{2})s(t)$ is expressed as $-\omega^{2}S(\omega)$ in the frequency domain, the high-frequency components, including the harmonics of the heartbeat, are emphasized; here, $\omega$ is the angular frequency and $S(\omega)$ represents the Fourier transform of $s(t)$.
@@ III-B2 Mel filters
$H_{\ell}(f)$ =$=\begin{cases}\vspace{2mm}\displaystyle\frac{2(f-f_{\ell})}{(f_{\ell+1}-f_{\ell})(f_{\ell+2}-f_{\ell})}&\displaystyle(f_{\ell}\leq f<f_{\ell+1}),\\ \vspace{2mm}\displaystyle\frac{2(f_{\ell+2}-f)}{(f_{\ell+2}-f_{\ell+1})(f_{\ell+2}-f_{\ell})}&\displaystyle(f_{\ell+1}\leq f<f_{\ell+2}),\\ 0&\text{otherwise}\end{cases}$ (11)
@@
$m_{\ell}=\tilde{m}\frac{\ell}{L+1}\log\left(1+f_{\mathrm{s}}/2\tilde{f}\right).$ (13)
@@
$S_{-\ell}$ $=\int_{0}^{T_{0}}\int_{-f_{\mathrm{s}}/2}^{0}|S(t,f)|H_{\ell}(-f)\,{\mathrm{d}}f{\mathrm{d}}t$ (16)
@@ III-B3 Discrete cosine transform
$C_{k}=$ $\frac{2}{L+\delta_{k0}}\sum_{\ell=0}^{L-1}\log S_{\ell}\cos\quantity{\frac{(2\ell+1)k\pi}{K}}$ (17)
for $k=0,1,\ldots,K-1$, and
$C_{k}=$ $\frac{2}{L+\delta_{k0}}\sum_{\ell=0}^{L-1}\log S_{-\ell}\cos\quantity{\frac{(2\ell+1)k\pi}{K}}$ (18)
for $k=-0,-1,-2,\ldots,-(K-1)$, where $\delta_{i,j}$ is the Kronecker delta, $K$ is the cepstrum dimension, and we set $K=64$. Note that $C_{+0}$ and $C_{-0}$ are also regarded as different variables here.
@@ IV-A Experiment 1: Five-Day Measurement of Six Participants
In this experiment, measurements were conducted over five days with six healthy adult participants. Each day, measurements were taken five times per participant in the morning and another five times per participant in the afternoon, with each session lasting for $60.0~{}\mathrm{s}$. Therefore, the total number of measurements performed was $(5+5)\times 5=50$ times per each person, giving a total of 300 measurements. All participants were in good health with no history of respiratory problems or heart disease. All measurements were conducted with each participant seated in front of the radar system at a distance of $1.5~{}\mathrm{m}$. The experimental setup is pictured in Fig. 6.
Figure 6: Experimental setup with the radar system and a seated participant.
The radar specifications are listed in Table II. We used a frequency-modulated continuous-wave (FMCW) radar system with a center frequency of $79.0~{}\mathrm{GHz}$ and a bandwidth of $3.6~{}\mathrm{GHz}$. The antenna array consists of three transmitting and four receiving elements, which form a MIMO array that can be approximated virtually as a 12-element linear array. The range and the direction of arrival of the echo from the target are identified by performing a Fourier transform in terms of the fast time  [54] and the virtual element number corresponding to the beamformer method [55].
TABLE II: SPECIFICATIONS OF THE RADAR SYSTEM
@@
Figure 7: Two-dimensional t-SNE visualizations of the feature vectors (a) $\bm{r}_{\mathrm{resp}}$, (b) $\bm{r}_{\mathrm{hb}}$, and (c) $\bm{r}_{\mathrm{prop}}$.
@@
The three feature vectors $\bm{r}_{\mathrm{resp}}$, $\bm{r}_{\mathrm{hb}}$, and $\bm{r}_{\mathrm{prop}}$ are input into the three classifiers to compare their identification accuracies. Here, because both SVM and $k$-NN are binary classifiers, the one-versus-rest method is used to perform multi-class classification [57]. In addition, the activation function used for the MLP output layer is the softmax function. The parameters for each classifier are optimized via a grid search. The optimization processes are performed for each of the three feature vectors $\bm{r}_{\mathrm{resp}}$, $\bm{r}_{\mathrm{hb}}$, and $\bm{r}_{\mathrm{prop}}$ individually. The parameters to be optimized and the values selected are given in Table III.
TABLE III: OPTIMIZED PARAMETERS FOR EACH METHOD
@@
Figure 8: Confusion matrices for each method. (a) SVM, (b) $k$-NN, (c) MLP.
@@
Table IV presents the F1 scores and the AUCs for each participant when $\bm{r}_{\mathrm{prop}}$ is used as the input, along with their respective macro-averages. From Table IV, we see that the SVM achieves the highest F1 scores and AUCs among the classifiers. The AUC exceeds a macro average of 0.99 for all classifiers, thus indicating that each classifier achieves comparable accuracy to the others. This indicates that the proposed features enable high-accuracy identification, regardless of the classifier that is used.
Figure 9: ROC curves for (a) SVM, (b) $k$-NN, and (c) MLP.
TABLE IV: F1 SCORE AND AUC FOR EACH PARTICIPANT USING $\bm{r}_{\mathrm{prop}}$ IN EXP. 1
@@ IV-B Experiment 2: Public Dataset of Thirty Participants
Next, we evaluated the accuracy of the proposed method using a public dataset provided by Schellenberger et al. [58], where they used a continuous-wave (CW) radar system with an operating frequency of $24~{}\mathrm{GHz}$ to collect echoes from a total of 30 individuals, comprising 14 male and 16 female participants. The echoes were recorded with a sampling frequency of $2.0\,\mathrm{kHz}$, which was then downsampled to $100\,\mathrm{Hz}$ to make the conditions consistent with those of experiment 1. The measurement time varied for each participant and ranged from approximately $600\,\mathrm{s}$ to $700\,\mathrm{s}$. In this study, we divided the data into segments with a length of $T_{0}$ and examined two cases, where $T_{0}=60.0\,\mathrm{s}$ and $T_{0}=5.0\,\mathrm{s}$; a feature vector was then generated from each segment,
Fig. 10 visualizes two-dimensional plots of the feature vector $\bm{r}_{\mathrm{hb}}$ as obtained from data segments with $T_{0}=60.0\,\mathrm{s}$ and $T_{0}=5.0\,\mathrm{s}$ when compressed using t-SNE. In both cases, it was observed that almost all plots were clearly separated into distinct clusters corresponding to the 30 participants. Next, the feature vector $\bm{r}_{\mathrm{hb}}$ obtained was input into the SVM and the identification accuracy was evaluated by performing a stratified $k$-fold cross-validation with $k=5$; the resulting confusion matrices are shown in Fig. 11. First, when $T_{0}=60.0\;\mathrm{s}$, 100% accuracy was achieved for all but four participants, and the overall accuracy was 98.69%, thus indicating that individual identification was achieved successfully. Next, when $T_{0}=5.0\,\mathrm{s}$, the overall accuracy realized was 99.39%, which also represents high-precision identification.
Figure 10: Two-dimensional plots of feature vectors $\bm{r}_{\mathrm{hb}}$ compressed by t-SNE when using the public dataset with a data length (a) $T_{0}=60.0$ s and (b) $T_{0}=5.0$ s.
Figure 11: Confusion matrices for experiment 2 with a data length (a) $T_{0}=60\;\mathrm{s}$ and (b) $T_{0}=5\;\mathrm{s}$.
In experiment 1, the accuracy of method B2 was only 93.33%, whereas in experiment 2, the same method B2 achieved accuracy of 98.69%, partly because the public dataset was collected within a day and was thus unaffected by daily variations in the physiological signals. We should also note that the accuracy obtained for $T_{0}=5.0\;\mathrm{s}$ was almost the same as the accuracy for $T_{0}=60.0\;\mathrm{s}$, which may indicate that long measurements are not necessary during individual identification if the daily variations in the physiological signals can be neglected.

### v2

@@ I Introduction
Recently, radar has been used widely for measurement of human bodies because both limb movements and physiological signals can be acquired in a noncontact manner and privacy concerns are reduced when compared with use of camera-based sensors [12, 13, 14]. Radar techniques have been used for respiration measurement  [15, 16, 17], heartbeat [18, 21, 23, 22, 19, 20] and also for individual identification applications [24, 25].
@@
In this paper, we propose an individual identification method that uses both respiratory and heartbeat features for the first time. A preprint of this manuscript has been posted online [53]. The contributions of this paper are listed as follows:
@@ II-A Waveform Feature Extraction Method
where $\lambda$ is the wavelength, $s(t)$ is the signal reflected from the target, $\mathrm{unwrap}\{\cdot\}$ is the phase unwrapping operator, and $\angle$ represents the phase of a complex number. When the target person is stationary, $d(t)$ mainly contains the displacement components that result from the person’s respiration and heartbeat.
@@
Example MRCW models are shown in Fig. 2 for $f=1$ Hz and $A=1$. As the figure shows, the waveform for the MRCW model is dependent on the parameters $\quantity(f,\beta_{1},\beta_{2},D)$. Hereafter, the parameters $\bm{p}=\quantity(f,\beta_{1},\beta_{2},D)^{\mathrm{T}}$ are used as the respiratory features (a superscript $\mathrm{T}$ signifies the transpose operator).
@@
where $T_{0}$ is the measurement time length, $\tau$ is the time shift and $\hat{d}_{\mathrm{MRCW}}(t)$ is the MRCW model that provides the best approximation of $d(t)$. In addition, $w(t)$ is a rectangular window with a length of 8 s. The parameters estimated in (5) are defined here as the estimated waveform features $\hat{\bm{p}}(t)=(\hat{f}(t),\hat{\beta}_{1}(t),\hat{\beta}_{2}(t),\hat{D}(t))^{\mathrm{T}}$ at a time $t$.
@@ II-B Features of the Respiratory Expiratory Plateau
$w_{\mathrm{E}}^{(t)}(t^{\prime})=\begin{cases}w(t^{\prime}-t)&\quantity(\hat{d}_{\mathrm{MRCW}}(t^{\prime})\geq\gamma\hat{A}(t)),\\ 0&(\text{otherwise}),\end{cases}$ (8)
where $\gamma=0.6$ is a threshold that was set empirically. In this paper, $\hat{c}_{2}(t)$ is used as a feature that represents the expiratory plateau’s shape. Fig. 4 shows the time interval for the expiratory plateau (gray area) and the corresponding estimated parabolic curve (red solid line).
@@ III-A High-Frequency Enhancement Using Differentiation
In this section, we propose a method to extract the features of the heartbeat component required for individual identification. Because the fundamental frequency component of the heartbeat is often masked by harmonic components of the respiration, it has proven to be advantageous to use the higher harmonics of the heartbeat components [18, 19, 20]. Therefore, our proposed method emphasizes the high-frequency components by using second-order differentiation, as in our previous work [23]. Because the second-order differentiation $s^{\prime\prime}(t)=({\mathrm{d}}^{2}/{\mathrm{d}}t^{2})s(t)$ is expressed as $-\omega^{2}S(\omega)$ in the frequency domain, the high-frequency components, including the harmonics of the heartbeat, are emphasized; here, $\omega$ is the angular frequency and $S(\omega)$ represents the Fourier transform of $s(t)$.
@@ III-B2 Mel filters
$H_{\ell}(f)$ =$=\begin{cases}\vspace{2mm}\displaystyle\frac{2(f-f_{\ell})}{(f_{\ell+1}-f_{\ell})(f_{\ell+2}-f_{\ell})}&\displaystyle(f_{\ell}\leq f<f_{\ell+1}),\\ \vspace{2mm}\displaystyle\frac{2(f_{\ell+2}-f)}{(f_{\ell+2}-f_{\ell+1})(f_{\ell+2}-f_{\ell})}&\displaystyle(f_{\ell+1}\leq f<f_{\ell+2}),\\ 0&(\text{otherwise})\end{cases}$ (11)
@@
$m_{\ell}=\tilde{m}\frac{\ell}{L+1}\log\left(1+\frac{f_{\mathrm{s}}}{2\tilde{f}}\right),$ (13)
@@
$S_{-\ell}$ $=\int_{0}^{T_{0}}\int_{-f_{\mathrm{s}}/2}^{0}|S(t,f)|H_{\ell}(-f)\,{\mathrm{d}}f{\mathrm{d}}t,$ (16)
@@ III-B3 Discrete cosine transform
$C_{k}=$ $\frac{2}{L+\delta_{k,0}}\sum_{\ell=0}^{L-1}\log S_{\ell}\cos\quantity{\frac{(2\ell+1)k\pi}{K}},$ (17)
$C_{-k}=$ $\frac{2}{L+\delta_{-k,0}}\sum_{\ell=0}^{L-1}\log S_{-\ell}\cos\quantity{\frac{-(2\ell+1)k\pi}{K}},$ (18)
for $k=0,1,\ldots,K-1$, where $\delta_{i,j}$ is the Kronecker delta, $K$ is the cepstrum dimension, and we set $K=64$. Note that $C_{+0}$ and $C_{-0}$ are also regarded as different variables here.
@@ IV-A Experiment 1: Five-Day Measurement of Six Participants
In this experiment, measurements were conducted over five days with six healthy adult participants. Each day, measurements were taken five times per participant in the morning and another five times per participant in the afternoon, with each session lasting for $60.0~{}\mathrm{s}$. Therefore, the total number of measurements performed was $(5+5)\times 5=50$ times per each person, giving a total of 300 measurements. In other words, a total of 300 samples were taken. All participants were in good health with no history of respiratory problems or heart disease. The physical characteristics of each of the participants are listed in Table II.
TABLE II: PHYSICAL CHARACTERISTICS OF EACH PARTICIPANT
Participant Age (yr) Height (m) Weight (kg)
1 26 1.67 58.9
2 22 1.76 69.7
3 23 1.75 75.2
4 23 1.80 62.0
5 23 1.70 68.0
6 23 1.69 54.0
All measurements were conducted with each participant seated in front of the radar system at a distance of $1.5~{}\mathrm{m}$. The experimental setup is pictured in Fig. 6, and a schematic that indicates the physical dimensions is shown in Fig. 6.
Figure 6: (a) Photograph and (b) schematic of the experimental setup showing the radar system and a seated participant.
The radar specifications are listed in Table III. We used a frequency-modulated continuous-wave (FMCW) radar system with a center frequency of $79.0~{}\mathrm{GHz}$ and a bandwidth of $3.6~{}\mathrm{GHz}$. The antenna array consists of three transmitting and four receiving elements, which form a MIMO array that can be approximated virtually as a 12-element linear array. The range and the direction of arrival of the echo from the target are identified by performing a Fourier transform in terms of the fast time  [57] and the virtual element number corresponding to the beamformer method [58].
TABLE III: SPECIFICATIONS OF THE RADAR SYSTEM
@@
Figure 7: Two-dimensional t-SNE visualizations of the feature vectors (a) $\bm{r}_{\mathrm{resp}}$, (b) $\bm{r}_{\mathrm{hb}}$, and (c) $\bm{r}_{\mathrm{prop}}$ for the 60 s data samples acquired in experiment 1.
@@
The three feature vectors $\bm{r}_{\mathrm{resp}}$, $\bm{r}_{\mathrm{hb}}$, and $\bm{r}_{\mathrm{prop}}$ are input into the three classifiers to compare their identification accuracies. Here, because both SVM and $k$-NN are binary classifiers, the one-versus-rest method is used to perform multi-class classification [60]. In addition, the activation function used for the MLP output layer is the softmax function. The parameters for each classifier are optimized via a grid search. The optimization processes are performed for each of the three feature vectors $\bm{r}_{\mathrm{resp}}$, $\bm{r}_{\mathrm{hb}}$, and $\bm{r}_{\mathrm{prop}}$ individually. The parameters to be optimized and the values selected are given in Table IV.
TABLE IV: OPTIMIZED PARAMETERS FOR EACH METHOD
@@
Figure 8: Confusion matrices for each method (a) SVM, (b) $k$-NN, and (c) MLP for the 60 s data samples acquired in experiment 1.
@@
Table V presents the F1 scores and the AUCs for each participant when $\bm{r}_{\mathrm{prop}}$ is used as the input, along with their respective macro-averages. From Table V, we see that the SVM achieves the highest F1 scores and AUCs among the classifiers. The AUC exceeds a macro average of 0.99 for all classifiers, thus indicating that each classifier achieves comparable accuracy to the others. This indicates that the proposed features enable high-accuracy identification, regardless of the classifier that is used.
Figure 9: ROC curves for (a) SVM, (b) $k$-NN, and (c) MLP for the 60 s data samples acquired in experiment 1.
TABLE V: F1 SCORE AND AUC FOR EACH PARTICIPANT USING $\bm{r}_{\mathrm{prop}}$ IN EXP. 1
@@
Next, to investigate the individual identification performance for short data samples, we divide each 60 s data sample into 12 short data samples with a length of 5.0 s, thus generating 3,600 samples in total. For these shorter samples, we evaluate the performance for the heartbeat feature $\bm{r}_{\mathrm{hb}}$ only (corresponding to methods B1, B2, and B3) because the data sample length of 5.0 s is not sufficient to capture respiratory features, which have a typical interval of between 3 s and 5 s.
The feature vector $\bm{r}_{\mathrm{hb}}$ for the short data samples is visualized in two-dimensional space using the t-SNE method in Fig. 10. Although the figure shows six clusters corresponding to the six participants, these clusters overlap severely when compared with those shown in Fig. 7; this is an indication of the performance degradation caused by the limited data sample length. The confusion matrices obtained using the stratified $k$-fold cross-validation approach are shown in Fig. 11 for $k=10$ for the 5 s data samples. The accuracies obtained for methods B1, B2, and B3 with the 5 s data samples were 85.64%, 83.11%, and 84.19%, respectively, showing that the performance has been reduced by 7.69, 9.22, and 3.81 points when compared with the accuracies of 93.33%, 92.33%, and 88.00% for B1, B2, and B3, respectively, obtained with the 60 s data samples. We also confirmed that the AUCs obtained for methods B1, B2, and B3 with the 5 s data samples were 0.979, 0.968, and 0.962, respectively, which were lower than the AUCs obtained with the 60 s data samples by 0.017, 0.024, and 0.022, respectively.
Figure 10: Two-dimensional t-SNE visualizations of the feature vector $\bm{r}_{\mathrm{hb}}$ for the 5 s data samples acquired in experiment 1.
Figure 11: Confusion matrices for each method: (a) SVM, (b) $k$-NN, and (c) MLP for the 5 s data samples acquired in experiment 1.
@@ IV-B Experiment 2: Public Dataset of Thirty Participants
Next, we evaluated the accuracy of the proposed method using a public dataset provided by Schellenberger et al. [61], where they used a continuous-wave (CW) radar system with an operating frequency of $24~{}\mathrm{GHz}$ to collect echoes from a total of 30 individuals, comprising 14 male and 16 female participants. The echoes were recorded with a sampling frequency of $2.0~{}\mathrm{kHz}$, which was then downsampled to $100~{}\mathrm{Hz}$ to make the conditions consistent with those of experiment 1. The measurement time varied for each participant and ranged from approximately $600~{}\mathrm{s}$ to $700~{}\mathrm{s}$. In this study, we divided the data into segments with a length of $T_{0}$ and examined two cases, where $T_{0}=60.0~{}\mathrm{s}$ and $T_{0}=5.0~{}\mathrm{s}$; therefore, totals of 306 and 3,795 samples were taken, respectively. A feature vector was then generated from each sample.
Fig. 12 visualizes two-dimensional plots of the feature vector $\bm{r}_{\mathrm{hb}}$ as obtained from data segments with $T_{0}=60.0~{}\mathrm{s}$ and $T_{0}=5.0~{}\mathrm{s}$ when compressed using t-SNE. In both cases, it was observed that almost all plots were clearly separated into distinct clusters corresponding to the 30 participants. Next, the feature vector $\bm{r}_{\mathrm{hb}}$ obtained was input into the SVM and the identification accuracy was evaluated by performing a stratified $k$-fold cross-validation with $k=5$; the resulting confusion matrices are shown in Fig. 13. First, when $T_{0}=60.0~{}\mathrm{s}$, 100% accuracy was achieved for all but four participants, and the overall accuracy was 98.69%, thus indicating that individual identification was achieved successfully. Next, when $T_{0}=5.0~{}\mathrm{s}$, the overall accuracy realized was 99.39%, which also represents high-precision identification.
Figure 12: Two-dimensional plots of feature vectors $\bm{r}_{\mathrm{hb}}$ compressed by t-SNE when using the public dataset with data lengths (a) $T_{0}=60.0$ s and (b) $T_{0}=5.0$ s in experiment 2.
Figure 13: Confusion matrices for experiment 2 with a data length (a) $T_{0}=60~{}\mathrm{s}$ and (b) $T_{0}=5~{}\mathrm{s}$.
In experiment 1, the accuracy of method B2 was only 93.33%, whereas in experiment 2, the same method B2 achieved accuracy of 98.69%, partly because the public dataset was collected within a day and was thus unaffected by daily variations in the physiological signals. We should also note that the accuracy obtained for $T_{0}=5.0~{}\mathrm{s}$ was almost the same as the accuracy for $T_{0}=60.0~{}\mathrm{s}$, which may indicate that long measurements are not necessary during individual identification if the daily variations in the physiological signals can be neglected.